# DEEP LEARNING-DRIVEN MALWARE CLASSIFICATION WITH API CALL SEQUENCE ANALYSIS AND CONCEPT DRIFT HANDLING

### v1

@@ 3 Related Work
Chen et al. [10] distinguished between feature-space drift and data-space drift in malware detectors, highlighting the predominant influence of data-space drift on model degradation over time. Their findings underscored the necessity for further exploration into the implications of feature-space updates, particularly in the context of Android malware datasets like AndroZoo and EMBER.
Jameel et al. [11] conducted a critical review of concept drift’s adverse effects on machine learning classification models. They proposed the ACNNELM model as optimal for Big Data stream classification but noted the absence of critical parameters for advanced ML models like deep learning. The review also highlighted the lack of a matrix model to measure adaptability factors, suggesting avenues for future research in model evaluation and optimization.
Hashmani et al. [12] presented a systematic literature review on concept drift evolution in machine learning approaches. Their comprehensive synthesis and categorization of existing research provided valuable insights into the state of the art in handling concept drift. However, the paper itself did not contribute new methodologies, serving primarily as a reference for researchers seeking to understand current trends and challenges in the field.
Lu et al. [2] conducted an extensive review focusing on concept drift in machine learning, covering detection, understanding, and adaptation strategies across numerous studies. While offering valuable insights into the breadth of research in this area, the paper did not introduce novel methodologies, functioning primarily as a compilation and analysis of existing approaches.
Farid et al. [13] proposed an adaptive ensemble classifier for mining concept drifting data streams. Their methodology addressed the challenge of concept drift by leveraging an ensemble approach, demonstrating promise for real-world applications with evolving data streams. However, the effectiveness of their approach was contingent upon the selection of appropriate base classifiers, highlighting a potential area for improvement in future research.
@@ 4 Data Collection and Preprocessing
In this section, we discuss the process of data collection from VirusShare [14] and VirusTotal [15].
@@
    Downloading Malware Hashes: The process starts with obtaining malware hashes from VirusShare [14], a platform known for its collection of malware samples.
@@
    Passing Hashes to VirusTotal: These hashes are then sent to VirusTotal [15] using an API key. VirusTotal analyzes files and URLs, stores the hashes, and signatures in their repository to detect malicious content.
@@ 5 Proposed Architecture
  2. 2.
    Calculation of Fitness Score: The fitness score [16] is calculated as the edit distance from the target string. The formula for calculating the fitness score for an individual in the population is:
    $\text{{Fitness}}(individual)=\sum_{i=1}^{n}1_{a_{i}\neq b_{i}}$ (1)
    Where:
    • $\text{{Fitness}}(individual)$ is the fitness score of the individual.
    • $n$ is the length of the target string.
    • $a_{i}$ is the $i$-th character of the target string.
    • $b_{i}$ is the $i$-th character of the individual.
  3. 3.
@@ 7 Comparison with Related work
García et al. [17] proposed a technique to assess the effectiveness of transfer learning (TL) methods for malware classification in the presence of concept drift, focusing on various time periods and learning scenarios. The study utilised five TL algorithms—TrAda, CORAL, DAE, DTS, and TIT— and evaluated their effectiveness in handling concept drift in malware classification. TrAda, CORAL, and DAE were identified as the most effective TL algorithms, consistently achieving Matthews correlation coefficients (MCC) greater than 0.9. Among the machine learning (ML) algorithms, Random Forest (RF) demonstrated competitive performance, especially in the inductive TL setting. The dataset used in the study was not explicitly disclosed by the authors, making the exact source of the malware and benign dataset unclear.
@@
1 García et al. [17] Transfer learning Not Disclosed 9196 APIs, signatures,
@@
2 Fernando et al. [18] Genetic Algorithms Elderan dataset 2720 API Calls
@@ 3 Proposed Work Genetic Algorithms VirusShare 22054 API Calls
Fernando and Komninos [18] introduced the feSAD ransomware detection framework, which leverages machine learning to adapt to concept drift. The framework aims at enhancing ransomware detection rates by calibrating drift thresholds and identifying abnormal drift samples. Compatible with various machine learning algorithms, it has demonstrated effectiveness in detecting ransomware amidst concept drift and can be tailored for different malware types. Experimental results, including 720 ransomware samples and 2000 benign samples from the Elderan dataset, show high detection rates and precision, especially with the Random Forest algorithm. Komninos highlighted the Random Forest’s stability in ransomware detection. However, they noted that in test batch 2, a high volume of ransomware samples caused abnormal concept drift, leading to reduced detection rates and statistical drift, suggesting system retraining to avoid detection decline.

### v2

@@ 3 Related Work
Chen et al. chen2023overkill distinguished between feature-space drift and data-space drift in malware detectors, highlighting the predominant influence of data-space drift on model degradation over time. Their findings underscored the necessity for further exploration into the implications of feature-space updates, particularly in the context of Android malware datasets like AndroZoo and EMBER.
Jameel et al. jameel2020critical conducted a critical review of concept drift’s adverse effects on machine learning classification models. They proposed the ACNNELM model as optimal for Big Data stream classification but noted the absence of critical parameters for advanced ML models like deep learning. The review also highlighted the lack of a matrix model to measure adaptability factors, suggesting avenues for future research in model evaluation and optimization.
Hashmani et al. hashmani2020concept presented a systematic literature review on concept drift evolution in machine learning approaches. Their comprehensive synthesis and categorization of existing research provided valuable insights into the state of the art in handling concept drift. However, the paper itself did not contribute new methodologies, serving primarily as a reference for researchers seeking to understand current trends and challenges in the field.
Lu et al. lu2018learning conducted an extensive review focusing on concept drift in machine learning, covering detection, understanding, and adaptation strategies across numerous studies. While offering valuable insights into the breadth of research in this area, the paper did not introduce novel methodologies, functioning primarily as a compilation and analysis of existing approaches.
Farid et al. farid2013adaptive proposed an adaptive ensemble classifier for mining concept drifting data streams. Their methodology addressed the challenge of concept drift by leveraging an ensemble approach, demonstrating promise for real-world applications with evolving data streams. However, the effectiveness of their approach was contingent upon the selection of appropriate base classifiers, highlighting a potential area for improvement in future research.
@@ 4 Data Collection and Preprocessing
In this section, we discuss the process of data collection from VirusShare virusshare and VirusTotal virustotal .
@@
    Downloading Malware Hashes: The process starts with obtaining malware hashes from VirusShare virusshare , a platform known for its collection of malware samples.
@@
    Passing Hashes to VirusTotal: These hashes are then sent to VirusTotal virustotal using an API key. VirusTotal analyzes files and URLs, stores the hashes, and signatures in their repository to detect malicious content.
@@ 5 Proposed Architecture
  2. 2.
    Calculation of Fitness Score: The fitness score kozeny2015genetic is calculated as the edit distance from the target string. The formula for calculating the fitness score for an individual in the population is:
    $\text{{Fitness}}(individual)=\sum_{i=1}^{n}1_{a_{i}\neq b_{i}}$ (1)
    Where:
    • $\text{{Fitness}}(individual)$ is the fitness score of the individual.
    • $n$ is the length of the target string.
    • $a_{i}$ is the $i$-th character of the target string.
    • $b_{i}$ is the $i$-th character of the individual.
@@ 7 Comparison with Related work
García et al. garcia2023effectiveness proposed a technique to assess the effectiveness of transfer learning (TL) methods for malware classification in the presence of concept drift, focusing on various time periods and learning scenarios. The study utilised five TL algorithms—TrAda, CORAL, DAE, DTS, and TIT— and evaluated their effectiveness in handling concept drift in malware classification. TrAda, CORAL, and DAE were identified as the most effective TL algorithms, consistently achieving Matthews correlation coefficients (MCC) greater than 0.9. Among the machine learning (ML) algorithms, Random Forest (RF) demonstrated competitive performance, especially in the inductive TL setting. The dataset used in the study was not explicitly disclosed by the authors, making the exact source of the malware and benign dataset unclear.
@@
1 García et al. garcia2023effectiveness Transfer learning Not Disclosed 9196 APIs, signatures,
@@
2 Fernando et al. fernando2024fesad Genetic Algorithms Elderan dataset 2720 API Calls
@@ 3 Proposed Work Genetic Algorithms VirusShare 22054 API Calls
Fernando and Komninos fernando2024fesad introduced the feSAD ransomware detection framework, which leverages machine learning to adapt to concept drift. The framework aims at enhancing ransomware detection rates by calibrating drift thresholds and identifying abnormal drift samples. Compatible with various machine learning algorithms, it has demonstrated effectiveness in detecting ransomware amidst concept drift and can be tailored for different malware types. Experimental results, including 720 ransomware samples and 2000 benign samples from the Elderan dataset, show high detection rates and precision, especially with the Random Forest algorithm. Komninos highlighted the Random Forest’s stability in ransomware detection. However, they noted that in test batch 2, a high volume of ransomware samples caused abnormal concept drift, leading to reduced detection rates and statistical drift, suggesting system retraining to avoid detection decline.